OpenStudy (anonymous):

find y'. y=tan^-1 x/2

zepdrix (zepdrix):

\[\large y=\arctan\left(\frac{x}{2}\right)\] Are you allowed to use the definition of the derivative of arctan? Or does your teacher want you to do the steps to work through it?

zepdrix (zepdrix):

\[\large \color{orangered}{y=\arctan x}\]\[\large \color{orangered}{y'=\frac{1}{1+x^2}}\]Are we allowed to use this? :)

OpenStudy (anonymous):

we have not learned the dervivative of arctan yet. So im guessing he want us to work out but i like that formula better so we can use it

zepdrix (zepdrix):

Hah ok let's use the formula, it will be rather quick :D If you wanna go through the steps after that we can do that. \[\large y=\arctan\left(\frac{x}{2}\right)\] \[\large y'=\frac{1}{1+\left(\dfrac{x}{2}\right)^2}\left(\frac{x}{2}\right)'\] Ok any questions about how I plugged that in? Instead of x our argument is x/2. So anywhere we would normally plug in x, we'll plug in x/2. We also need to apply the chain rule. The prime on the outside of that bracket lets us know that we still need to take the derivative of that part.

OpenStudy (anonymous):

the derivative of x/2 is 1/2 right? and then what do i do after that?

zepdrix (zepdrix):

Yes good good c: sorry got distracted lol

OpenStudy (anonymous):

or is it 1-x/2

zepdrix (zepdrix):

Then from there, just simplify if you want :D\[\large y'=\frac{\left(\dfrac{1}{2}\right)}{1+\left(\dfrac{x}{2}\right)^2}\]If we multiply the top and bottom by 4 we get,\[\large y'=\frac{2}{4+x^2}\]Which looks a little bit nicer.

zepdrix (zepdrix):

1-x/2? :o

OpenStudy (anonymous):

oh okay i got it. and sorry i sometimes over analyze stuff. lol

zepdrix (zepdrix):

heh ^^